# The Wavefront Set: Bounds for the Langlands parameter

### v2

@@ Abstract.
We are interested in the interplay between the Langlands parameter of a smooth irreducible representation of the reductive $p$-adic group and the geometric wavefront set of the representation, an invariant that appears naturally in the local character expansion of Howe [Ho] and Harish-Chandra [HC].
@@ 1. Introduction: a conjecture
Let ${\mathsf{k}}$ be a nonarchimedean field with residue field $\mathbb{F}_{q}$. Let $\bar{\mathsf{k}}$ be a separable closure of ${\mathsf{k}}$ and let $K\subset\bar{\mathsf{k}}$ denote the maximal unramified extension. Let $W_{\mathsf{k}}$ be the Weil group with inertia subgroup $I_{\mathsf{k}}$ and norm ||||||~{}||| | | |. Let $W^{\prime}_{\mathsf{k}}$ be the Weil-Deligne group, i.e., $W^{\prime}_{\mathsf{k}}=W_{\mathsf{k}}\ltimes{\mathbb{C}}$, where the action is defined by $w\cdot z=||w||z$, $w\in W_{\mathsf{k}}$, $z\in{\mathbb{C}}$. Let $\mathsf{Fr}$ denote the geometric Frobenius element, which generates $W_{\mathsf{k}}/I_{\mathsf{k}}$.
Let ${\mathbf{G}}$ be a reductive ${\mathsf{k}}$-group that splits over an unramified extension. The pure inner forms of ${\mathbf{G}}$, in the sense of [Vo] are parametrized by the first Galois cohomology group $H^{1}({\mathsf{k}},{\mathbf{G}})$. For every $\omega\in H^{1}({\mathsf{k}},{\mathbf{G}})$, let ${\mathbf{G}}^{\omega}({\mathsf{k}})$ denote the corresponding pure inner form. Let $G^{\vee}$ be the complex dual group to ${\mathbf{G}}$. Let ${}^{L}\!G=G^{\vee}\rtimes W_{\mathsf{k}}$ be the L-group.
@@ 1.1. Langlands parameters
We briefly recall the conjectural Langlands parametrization of the irreducible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representations, in the formulation of [Vo], generalizing the Kazhdan-Lusztig geometric parametrization of representations with Iwahori-fixed vectors. A Langlands parameter is a group homomorphism
@@
where $\varphi$ is a Langlands parameter and $\phi$ is an irreducible representation of $A_{\varphi}$. The pairs $(\varphi,\phi)$ are called enhanced Langlands parameters. Let $\Pi(\varphi)$ denote the Langlands packet, i.e., the set of representations parametrized by a fixed ($G^{\vee}$-orbit of) $\varphi$. We need the geometric reformulation of this correspondence.
Fix a group homomorphism $\lambda:I_{\mathsf{k}}\to{}^{L}\!G$ with finite image. The enhanced Langlands parameters for which $\varphi|_{I_{\mathsf{k}}}=\lambda$ admit the following interpretation. Let $H^{\vee}={G^{\vee}}^{\lambda(I_{\mathsf{k}})}$, a possibly disconnected reductive group, and ${\mathfrak{h}}^{\vee}={\mathfrak{g}^{\vee}}^{\lambda(I_{\mathsf{k}})}$, a reductive subalgebra of $\mathfrak{g}^{\vee}$, denote the fixed points.
Let $s=\varphi(\mathsf{Fr})\in{}^{L}\!G$ be a semisimple element. Since $\mathsf{Fr}$ normalizes $I_{\mathsf{k}}$, it follows that $s$ normalizes $H^{\vee}$ (and ${\mathfrak{h}}^{\vee}$). Let ${\mathfrak{h}}^{\vee}_{q}$ denote the $q$-eigenspace of $\mathrm{Ad}(s)$. Let $H^{\vee}(s)$ be the centralizer of $s$ in $H^{\vee}$. Then $H^{\vee}(s)$ acts on ${\mathfrak{h}}^{\vee}_{q}$ with finitely many orbits. There is a one-to-one correspondence between $G^{\vee}$-orbits of enhanced Langlands parameters $(\varphi,\phi)$ such that $\varphi(\mathsf{Fr})=s$ and $\varphi|_{I_{\mathsf{k}}}=\lambda$, and pairs $({{\mathcal{O}}}^{\vee},\mathcal{L}^{\vee})$ where
@@
Denote by $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})=\Pi(\varphi)$ the corresponding L-packet, each of whose member $\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ should be parametrized by the equivariant local systems $\mathcal{L}^{\vee}$ on ${{\mathcal{O}}}^{\vee}$. A desideratum is that $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$ consists of tempered representations if and only if ${{\mathcal{O}}}^{\vee}$ is dense in ${\mathfrak{h}}^{\vee}_{q}$.
1.2. Aubert-Zelevinsky involution
Let ${\mathsf{AZ}}$ denote the normalized Aubert-Zelevinsky involution [Au, Definition 1.5, Corollary 3.9] on $\mathsf{Irr}{\mathbf{G}}^{\omega}({\mathsf{k}})$ for every form ${\mathbf{G}}^{\omega}({\mathsf{k}})$. If $\varphi_{\pi}$ is the Langlands parameter of $\pi$, then it is expected that $\varphi_{\pi}(W_{\mathsf{k}})$ and $\varphi_{{\mathsf{AZ}}(\pi)}(W_{\mathsf{k}})$ are $G^{\vee}$ conjugate. Moreover, if $\pi=\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$, it is expected that the Langlands parameter of ${\mathsf{AZ}}(\pi)$ is related to that of $\pi$ via the Fourier transform [KS, §3.7] on the appropriate space of perverse sheaves of Langlands parameters, see also [EM, §3]. This is known to hold for all representations with Iwahori-fixed vectors by [EM, Theorem 0.1], and more generally it is expected to hold for unipotent representations in the sense of Lusztig.
More precisely, suppose that $\pi$ has Iwahori-fixed vectors, so $\lambda(I_{\mathsf{k}})=1$ and that $G^{\vee}$ is simply connected. The geometric space of parameters consists of irreducible $G^{\vee}(s)$-equivariant perverse sheaves on the complex vector space $\mathfrak{g}^{\vee}_{q}$. The connected reductive group $G^{\vee}(s)$ acts with finitely many orbits on $\mathfrak{g}^{\vee}_{q}$. This implies that $G^{\vee}(s)$ also acts with finitely many orbits on the dual vector space $(\mathfrak{g}^{\vee}_{q})^{*}$. Identify $(\mathfrak{g}^{\vee}_{q})^{*}$ with $\mathfrak{g}^{\vee}_{q^{-1}}$ by using a fixed nondegenerate invariant symmetric form on $\mathfrak{g}^{\vee}$. Pyasetsky’s duality [Ze] says that there exists a one-to-one correspondence between the orbits in $\mathfrak{g}^{\vee}_{q}$ and in $\mathfrak{g}^{\vee}_{q^{-1}}$, ${{\mathcal{O}}}^{\vee}\to P({{\mathcal{O}}}^{\vee})$, given as follows. For each ${{\mathcal{O}}}^{\vee}\subset\mathfrak{g}^{\vee}_{q}$, $P({{\mathcal{O}}}^{\vee})\subset\mathfrak{g}^{\vee}_{q^{-1}}$ is the unique orbit such that the conormal bundles have the same closures
@@ 1.2. Aubert-Zelevinsky involution
Let $\operatorname{FT}$ denote the Fourier transform from $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q}$ to $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q^{-1}}$, [EM, §3.2]. Let $\mathsf{IC}({{\mathcal{O}}},{{\mathcal{L}}})$ denote the simple perverse sheaf given by the intersection cohomology complex defined by the equivariant irreducible local system ${{\mathcal{L}}}$ supported on ${{\mathcal{O}}}$. Then $\operatorname{FT}(\mathsf{IC}({{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))=\mathsf{IC}({{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}})$ for an orbit ${{\mathcal{O}}}^{\vee}_{\mathsf{FT}}$ in $\mathfrak{g}^{\vee}_{q^{-1}}$ and an equivariant local system ${{\mathcal{L}}}^{\vee}_{\mathsf{FT}}$. If $\mathfrak{g}^{\vee}=\mathfrak{gl}(n)$, so that all local systems that appear are trivial, [EM, Proposition 7.2] shows that
@@
${\mathsf{AZ}}(\pi(s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))^{*}=\pi(s^{-1},{{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}}),$ (1.2)
where * in the left hand side denotes the contragredient. In our conjectures below, we will be interested in the nilpotent part of the Langlands parameter of ${\mathsf{AZ}}(\pi)$, that is in $\mathbb{O}^{\vee}_{\operatorname{FT}}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$
@@ 1.3. Spaltenstein and Lusztig-Sommers dualities
Let $G$ denote a complex group with maximal torus $T$ with dual $G^{\vee}$ and dual maximal torus $T^{\vee}$. Let $W$ denote the Weyl group of $(G,T)$ and $W^{\vee}$ the Weyl group of $(G^{\vee},T^{\vee})$. Of course, $W\cong W^{\vee}$. Let ${{\mathcal{U}}}_{G}$ and ${{\mathcal{U}}}_{G^{\vee}}$ denote the sets of unipotent elements in $G$ and $G^{\vee}$ respectively. Let $G\backslash{{\mathcal{U}}}_{G}$, $G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}$ denote the corresponding sets of conjugacy classes.
Let $d:G\backslash{{\mathcal{U}}}_{G}\to G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}$ and $d^{\vee}:G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}\to G\backslash{{\mathcal{U}}}_{G}$ denote the duality maps defined by Spaltenstein [Sp]. The images of $d,d^{\vee}$ consist of the set of special unipotent conjugacy classes.
Let $\mathfrak{g}$ be the Lie algebra of $G$. We will also make use of the duality map $d_{S}:G\backslash G\to G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}$ from conjugacy classes of $G$ to unipotent conjugacy classes of $G^{\vee}$, defined by Lusztig [Lu1, §13.3] in the case of special conjugacy classes and generalized by Sommers [So]. Suppose $x\in G$ has a Jordan decomposition $x=su$ and assume $s\in T$. Let $W(s)$ denote the centralizer of $s$ in $W$, this is the Weyl group of the centralizer $G(s)$. Let $E^{W(s)}_{u,1}$ denote the Springer $W(s)$-representation attached to the trivial local system for the unipotent class of $u$ in $G(s)$. Consider the induced representation $\operatorname{Ind}_{W(s)}^{W}(E^{W(s)}_{u,1})$. There exists a unique irreducible $W$-representation $\hat{E}$ in this induced representation such that $b_{\hat{E}}=b_{E^{W(s)}_{u,1}}$, where $b$ denotes the lowest harmonic degree. Regard $\hat{E}$ as an irreducible representation of $W^{\vee}$. By Springer’s correspondence for $G^{\vee}$, there exists a unique unipotent conjugacy class $u^{\vee}$ in $G^{\vee}$ such that $\hat{E}$ is realized in the top cohomology of the Springer fiber for $u^{\vee}$. Define $d_{S}(x)=u^{\vee}$. We recall that
@@ 1.4. Unramified wavefront set
$({{\mathcal{O}}},C):\ {{\mathcal{O}}}\text{ nilpotent ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit in }\mathfrak{g}(\bar{\mathsf{k}}),\ C\text{ conjugacy class in }A({{\mathcal{O}}}),$
where $A({{\mathcal{O}}})$ is a group of components of the centralizer $C_{{\mathbf{G}}(\bar{\mathsf{k}}}(u)$ in ${\mathbf{G}}(\bar{\mathsf{k}})$ of a representative $u$ of ${{\mathcal{O}}}$. By abuse of notation, we think of ${}^{K}\!\mathsf{WF}(\pi)$ as a collection of such pairs $({{\mathcal{O}}},C)$. For each ${{\mathcal{O}}}$ choose a representative $u\in{{\mathcal{O}}}$. For each $C$, we may choose a semisimple representative $\tau$ in $C_{{\mathbf{G}}(\bar{\mathsf{k}})}(u)$. Choosing an isomorphism $\bar{k}\cong\mathbb{C}$, we may identify ${\mathbf{G}}(\bar{\mathsf{k}})$ with the complex group $G={\mathbf{G}}(\mathbb{C})$ with the same root datum, and think of the duality maps in section 1.3 as maps between ${\mathbf{G}}(\bar{\mathsf{k}})$ and $G^{\vee}$. Write
@@
$({{\mathcal{O}}},\bar{C}):\ {{\mathcal{O}}}\text{ nilpotent $G$-orbit in }\mathfrak{g},\ \bar{C}\text{ conjugacy class in }\bar{A}({{\mathcal{O}}}),$

### v3

@@ Abstract.
For an irreducible smooth representation of a connected reductive $p$-adic group, two important associated invariants are the wavefront set and the (partly conjectural) Langlands parameter. While a wavefront set consists of $p$-adic nilpotent orbits, one constituent of the Langlands parameter is a complex nilpotent orbit in the dual Lie algebra. For unipotent representations in the sense of Lusztig, the corresponding nilpotent orbits on the two sides are related via the Lusztig–Spaltenstein duality, [CMBO1, CMBO2]. In this paper, we formulate a general upper-bound conjecture and several variants relating the nilpotent orbits that appear in the wavefront set and in the Langlands parameter. We also verify these expectations in some cases, including the depth-zero supercuspidal representations of classical groups and all the irreducible representations of $G_{2}$.
2020 Mathematics Subject Classification:
22E50, 22E35
@@ 1. Introduction: a conjecture
[…]
Let $\mathbf{G}$} be a connected reductive $\mathsf{k}$}-group that splits over an unramified extension. The pure inner forms of $\mathbf{G}$}, in the sense of [Vo, Definition 2.6] are parametrized by the first Galois cohomology group $H^{1}({\mathsf{k}},{\mathbf{G}})$. For every $\omega\in H^{1}({\mathsf{k}},{\mathbf{G}})$, let ${\mathbf{G}}^{\omega}({\mathsf{k}})$ denote the corresponding pure inner form. Let $G^{\vee}$ be the complex dual group to $\mathbf{G}$}. Let $^{L}\!G=G^{\vee}\rtimes W_{\mathsf{k}}$ be the $L$-group.
@@ 1.1. Langlands parameters
We briefly recall the conjectural Langlands parametrization of the irreducible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representations, in the formulation of [Vo, §4], generalizing the Kazhdan–Lusztig geometric parametrization of representations with Iwahori-fixed vectors [KL]. A Langlands parameter is a group homomorphism
@@
where $\varphi$ is a Langlands parameter and $\phi$ is an irreducible representation of $A_{\varphi}$. The pairs $(\varphi,\phi)$ are called enhanced Langlands parameters. Let $\Pi(\varphi)$ denote the Langlands packet, i.e., the set of representations parametrized by a fixed ($G^{\vee}$-orbit of) $\varphi$. We need a geometric reformulation of this correspondence.
Fix a group homomorphism $\lambda:I_{\mathsf{k}}\to{}^{L}\!G$ with finite image. The set of enhanced Langlands parameters for which $\varphi|_{I_{\mathsf{k}}}=\lambda$ admits the following interpretation. Let $H^{\vee}={G^{\vee}}^{\lambda(I_{\mathsf{k}})}$, a possibly disconnected reductive group, and ${\mathfrak{h}}^{\vee}={\mathfrak{g}^{\vee}}^{\lambda(I_{\mathsf{k}})}$, a reductive subalgebra of $\mathfrak{g}^{\vee}$, denote the fixed points.
Let $s$ be $\varphi(\mathsf{Fr})\in{}^{L}\!G$, a semisimple element. Since $\mathsf{Fr}$ normalizes $I_{\mathsf{k}}$, it follows that $s$ normalizes $H^{\vee}$ (and $\mathfrak{h}$$^{\vee}$). Let ${\mathfrak{h}}^{\vee}_{q}$ denote the $q$-eigenspace of $\mathrm{Ad}(s)$. Let $H^{\vee}(s)$ be the centralizer of $s$ in $H^{\vee}$. Then $H^{\vee}(s)$ acts on ${\mathfrak{h}}^{\vee}_{q}$ with finitely many orbits. There is a one-to-one correspondence between $G^{\vee}$-orbits of enhanced Langlands parameters $(\varphi,\phi)$ such that $\varphi(\mathsf{Fr})=s$ and $\varphi|_{I_{\mathsf{k}}}=\lambda$, and pairs $({{\mathcal{O}}}^{\vee},\mathcal{L}^{\vee})$ where
@@
Denote by $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})=\Pi(\varphi)$ the corresponding $L$-packet, each of whose member $\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ should be parametrized by the equivariant local systems $\mathcal{L}^{\vee}$ on $\mathcal{O}$$^{\vee}$. A desideratum is that if $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$ consists of tempered representations then $\mathcal{O}$$^{\vee}$ is dense in ${\mathfrak{h}}^{\vee}_{q}$.
1.2. Aubert–Zelevinsky involution
Let $\mathsf{AZ}$} denote the normalized Aubert–Zelevinsky involution [Au, Definition 1.5, Corollary 3.9] on $\mathsf{Irr}\,{\mathbf{G}}^{\omega}({\mathsf{k}})$ for every form ${\mathbf{G}}^{\omega}({\mathsf{k}})$. If $\varphi_{\pi}$ is the Langlands parameter of $\pi$, then it is expected that $\varphi_{\pi}(W_{\mathsf{k}})$ and $\varphi_{{\mathsf{AZ}}(\pi)}(W_{\mathsf{k}})$ are $G^{\vee}$ conjugate. Assuming this, denote the enhanced Langlands parameter of ${\mathsf{AZ}}(\pi)$ as follows
${\mathsf{AZ}}(\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))=\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{{\mathsf{AZ}}},{{\mathcal{L}}}^{\vee}_{{\mathsf{AZ}}}).$ (1.2)
It is expected that the enhanced Langlands parameter of ${\mathsf{AZ}}(\pi)$ is related to that of $\pi$ via the Fourier transform [KS, §3.7] on the appropriate space of perverse sheaves of Langlands parameters, see also [EM, §3]. This is known to hold for all representations with Iwahori-fixed vectors by [EM, Theorem 0.1].
More precisely, suppose that $\pi$ has Iwahori-fixed vectors, so $\lambda(I_{\mathsf{k}})=1$ and that $G^{\vee}$ is simply connected. The geometric space of parameters consists of irreducible $G^{\vee}(s)$-equivariant perverse sheaves on the complex vector space $\mathfrak{g}^{\vee}_{q}$. The connected reductive group $G^{\vee}(s)$ acts with finitely many orbits on $\mathfrak{g}^{\vee}_{q}$. This implies that $G^{\vee}(s)$ also acts with finitely many orbits on the dual vector space $(\mathfrak{g}^{\vee}_{q})^{*}$. Identify $(\mathfrak{g}^{\vee}_{q})^{*}$ with $\mathfrak{g}^{\vee}_{q^{-1}}$ by using a fixed nondegenerate invariant symmetric form on $\mathfrak{g}^{\vee}$. Pyasetsky’s duality [Ze] says that there exists a one-to-one correspondence between the orbits in $\mathfrak{g}^{\vee}_{q}$ and in $\mathfrak{g}^{\vee}_{q^{-1}}$, ${{\mathcal{O}}}^{\vee}\to P({{\mathcal{O}}}^{\vee})$, given as follows: for each ${{\mathcal{O}}}^{\vee}\subset\mathfrak{g}^{\vee}_{q}$, $P({{\mathcal{O}}}^{\vee})\subset\mathfrak{g}^{\vee}_{q^{-1}}$ is the unique orbit such that the conormal bundles have the same closures
@@ 1.2. Aubert–Zelevinsky involution
Let $\operatorname{FT}$ denote the Fourier transform from $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q}$ to $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q^{-1}}$, [EM, §3.2]. Let $\mathsf{IC}({{\mathcal{O}}},{{\mathcal{L}}})$ denote the simple perverse sheaf given by the intersection cohomology complex defined by the equivariant irreducible local system $\mathcal{L}$}} supported on $\mathcal{O}$}}. Then $\operatorname{FT}(\mathsf{IC}({{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))=\mathsf{IC}({{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}})$ for an orbit ${{\mathcal{O}}}^{\vee}_{\mathsf{FT}}$ in $\mathfrak{g}^{\vee}_{q^{-1}}$ and an equivariant local system ${{\mathcal{L}}}^{\vee}_{\mathsf{FT}}$. We emphasize that ${{\mathcal{O}}}^{\vee}_{\mathsf{FT}}$ depends on $\mathcal{L}$$^{\vee}$ as well, not just on $\mathcal{O}$$^{\vee}$. If $\mathfrak{g}^{\vee}=\mathfrak{gl}(n)$, so that all local systems that appear are trivial, [EM, Proposition 7.2] shows that
@@
${\mathsf{AZ}}(\pi(1,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))^{*}=\pi(1,s^{-1},{{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}}),$ (1.3)
where * in the left hand side denotes the contragredient. In terms of the previous notation, this means, in particular, that for Iwahori-spherical representations, $G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{{\mathsf{AZ}}}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}$. In our conjectures below, we will be interested in the nilpotent part of the Langlands parameter of ${\mathsf{AZ}}(\pi)$, that is, in $\mathbb{O}^{\vee}_{{\mathsf{AZ}}}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{AZ}}.$
@@ 1.3. Spaltenstein and Lusztig-Sommers dualities
[…]
Let $d:G\backslash{{\mathcal{U}}}_{G}\to G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}$ and $d^{\vee}:G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}\to G\backslash{{\mathcal{U}}}_{G}$ denote the duality maps defined by Spaltenstein [Sp, §10]. The images of $d,d^{\vee}$ consist of the set of special unipotent conjugacy classes.
[…]
@@ 1.4. Unramified wavefront set
[…]
where $A({{\mathcal{O}}})$ is a group of components of the centralizer $C_{{\mathbf{G}}(\bar{\mathsf{k}})}(u)$ in ${\mathbf{G}}(\bar{\mathsf{k}})$ of a representative $u$ of $\mathcal{O}$}}. By abuse of notation, we think of $^{K}\mathsf{WF}(\pi)$ as a collection of such pairs $({{\mathcal{O}}},C)$. For each $\mathcal{O}$}} choose a representative $u\in{{\mathcal{O}}}$. For each $C$, we may choose a semisimple representative $\tau$ in $C_{{\mathbf{G}}(\bar{\mathsf{k}})}(u)$. Choosing an isomorphism $\bar{\mathsf{k}}\cong\mathbb{C}$, we may identify ${\mathbf{G}}(\bar{\mathsf{k}})$ with the complex group $G={\mathbf{G}}(\mathbb{C})$ with the same root datum, and think of the duality maps in section 1.3 as maps between ${\mathbf{G}}(\bar{\mathsf{k}})$ and $G^{\vee}$. Write
@@
[…]